do the functions y1 (t)= t and y 2 ( t ) =1\ t form a fundamental sets of solutions of the equation 2t 2 y" + 3t y' - y=0, on the interval 0 <t< ∞ ? justify your answer.


1
Expert's answer
2021-05-28T10:21:18-0400

Let "y=t^\alpha". Then


"y'=\alpha t^{\alpha-1}"

"y''=\alpha(\alpha-1) t^{\alpha-2}"


Substitute


"2t^2\alpha(\alpha-1) t^{\alpha-2}+3t\alpha t^{\alpha-1}-t^\alpha=0"

"(2\alpha^2-2\alpha+3\alpha-1)t^\alpha=0"

"(2\alpha^2+\alpha-1)t^\alpha=0"

This is true for "0<t<\infin," if


"2\alpha^2+\alpha-1=0"

"2\alpha^2+2\alpha-\alpha-1=0"

"2\alpha(\alpha +1)-(\alpha +1)=0"

"2(\alpha +1)(\alpha-\dfrac{1}{2})=0"

"\alpha_1=\dfrac{1}{2}, \alpha_2=-1"

"y_1=t^{1/2}=\sqrt{t}"

"y_2=t^{-1}=\dfrac{1}{t}"

Hence "y_1(t)=\sqrt{t}" and "y_2(t)=\dfrac{1}{t}" form a fundamental sets of solutions of the equation 2t y" + 3t y' - y=0, on the interval 0 <t< ∞ .